40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 22 16:55:16 2012, Jamie, Update, SUS, wacky state of SUS input matrices 
    Reply  Thu Feb 23 02:20:07 2012, Zach, Update, SUS, wacky state of SUS input matrices 
       Reply  Wed Mar 7 21:13:30 2012, rana, Update, DAQ, NDS2 
    Reply  Fri Feb 24 04:12:44 2012, kiwamu, Update, SUS, freeswing test 
       Reply  Fri Feb 24 08:06:52 2012, steve, Update, SUS, sus restored freePRM.pngnoisyPRM.png
Message ID: 6312     Entry time: Fri Feb 24 08:06:52 2012     In reply to: 6311
Author: steve 
Type: Update 
Category: SUS 
Subject: sus restored 

Quote:
The following optics were kicked:
MC1 MC2 MC3 ETMX ETMY ITMX ITMY PRM SRM BS
Fri Feb 24 04:11:15 PST 2012
1014120690
 
Steve (or anyone), can you restore the watchdogs when you come to the lab in the morning ?

Quote from #6305

Kiwamu (or whoever is here last tonight): please run the free-swing/kick script (/opt/rtcds/caltech/c1/scripts/SUS/freeswing) before you leave, and I'll check the matrices and update the suspensions tomorrow morning.

 All suspentions were restored and MC locked. PRM side osem  RMS motion was high.

Atm2, Why the PRM is 2x as noisy as the SRM ?

Attachment 1: freePRM.png  46 kB  | Show | Hide all | Show all
Attachment 2: noisyPRM.png  36 kB  Uploaded Fri Feb 24 09:52:56 2012  | Hide | Hide all | Show all
noisyPRM.png
ELOG V3.1.3-